# UniGraph: Learning a Unified Cross-Domain Foundation Model for Text-Attributed Graphs

### v2

@@ Abstract.
Graph Neural Networks, Self-Supervised Learning, Graph Pre-Training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Neural networksccs: Information systems Social networks
@@ 4.2. Graph Siamese Masked Autoencoders on TAGs
The existing graph self-supervised learning methods (Veličković et al., 2018b; Kipf and Welling, 2016b; Thakoor et al., 2021; Hou et al., 2022) adopt GNNs as the backbone networks and use pre-processed vector features as input. In this part, we propose an end-to-end self-supervised learning method on TAGs. We cascade a pre-trained language model (LM) and a GNN as the backbone network. In this paper, we adopt DeBERTa-base (He et al., 2020) as our LM and GAT (Veličković et al., 2018a) as our GNN. Inspired by successful self-supervised learning techniques such as masked modeling (Kenton and Toutanova, 2019; He et al., 2022) and siamese networks (Wu et al., 2018), we designed Graph Siamese Masked Autoencoders, enabling large-scale self-supervised pre-training on TAGs.
@@
4.3. Graph Instruction Tuning
Instruction prompts. For instruction tuning, we design prompt templates that include graph embeddings, graph structure, and natural language instructions, as summarized in Appendix G.
@@ 4.3. Graph Instruction Tuning
    Due to the difference between the pre-trained datasets used in their original paper and ours, we re-run the experiments using their open-source code. Comparisons between UniGraph and their method using their original pre-trained datasets can be found in Appendix J.
@@
5. Experiments
5.1. Self-Supervised Representation Learning
Setup. We adopt the most commonly used linear probing protocol to evaluate the representation learning ability of self-supervised pre-trained models on unseen datasets. We train a linear classifier on top of the embeddings from a frozen model. Our model and all self-supervised learning baseline methods are first pre-trained on the large-scale citation network ogbn-Papers100M (Hu et al., 2020). Then, we evaluate on nine graphs from five domains with different tasks. Detailed settings and hyper-parameters can be found in Appendix C.
5.2. Few-Shot Transfer
5.3. Zero-Shot Transfer
Setup. For zero-shot transfer, we mainly compare with general-purpose LLMs such as Llama (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) and vicuna-7B-v1.5 (Chiang et al., 2023), as well as graph-based LLM methods [^1] Due to the use of different pre-training datasets, comparisons with other graph zero-shot learning baselines, such as LLaGA (Chen et al., 2024) and OpenGraph (Xia et al., 2024), can be found in Appendix J. GraphGPT (Tang et al., 2023), ZeroG (Li et al., 2024) and G2P2 (Wen and Fang, 2023). Also, we compare with OFA mentioned in section 5.2. Consistent with OFA, we select ogbn-Arxiv and FB15K237 as instruction tuning datasets. Note that there is no overlap between their label categories and those of the downstream datasets. We report the average accuracy on the original test set with 3 random initializations for each downstream dataset.
@@ Appendix A Datasets
Appendix B The Acquisition of Text-Attributed Graphs
TAGs are not a newly emerged, narrowly applied research topic. The benchmark datasets widely used in the field of graph learning (such as Cora, PubMed, OGB, etc.) are inherently TAGs, but researchers have pre-encoded them into vector features using methods like word2vec. This is also a major motivation for us to attempt to train a graph foundation model from TAGs: a large amount of graph data is collected from the web, where text and graphs naturally have a close relationship.
Automatic Textual Description Generation. Using natural language to replace vector encoding for representing graph semantic features, converting non-TAGs to TAGs can be an effective method. Although our model can only be used on TAGs, in the paper we also explored the effects of converting non-TAGs, such as molecular graphs, into TAGs. We further conduct experiments to demonstrate that this method of constructing features is comparable, or even superior, to standard vector features. In Table 11, “Standard” refers to standard 9-dimensional node features, containing atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring or not. “Text” represents the corresponding textual features we construct, while “text+word2vec” indicates that we use word2vec to encode these texts into vectors.
Table 11. Performance comparison of different feature construction methods on HIV and PCBA datasets.
Features Method HIV PCBA
Standard GAT 70.89$\pm$1.23 56.11$\pm$1.23
Text+word2vec GAT 71.12$\pm$1.34 56.24$\pm$1.01
Text UniGraph 71.23$\pm$1.93 57.67$\pm$0.85
Similarly, for graph data without features, the traditional approach is to construct vector features from each node’s structural features (such as node degree, etc.). We can also use natural language to describe these semantics, thereby converting them into TAGs.
Appendix C Implementation Notes
@@
Table 12. Pre-training hyper-parameters for our framework on ogbn-Papers100M.
mask rate hidden_size lr weight_decay dropout optimizer num_epochs num_gnn_layers ppr topk ema decay coefficient $\lambda$
0.75 768 2e-5 0.001 0.2 adamw 1 3 128 0.996 0.1
Table 13. Graph instruction tuning hyper-parameters for our framework on ogbn-Arxiv and FB15K237.
LLM hidden_size lr weight_decay dropout optimizer num_epochs warmup_ratio clip_grad_norm batch_size max_text_length
Llama 4096 8e-5 0.0 0.0 adamw 2 0.05 1.0 4 1024
Table 14. Analysis of LLMs Choices. We report accuracy (%) for all datasets.
Cora PubMed Products1 Wiki-CS WN18RR
Llama2-13B 37.34 62.43 50.11 24.23 37.98 44.24 25.32 36.56 17.35 18.93 34.12
OFA 24.01 56.92 54.01 - - - - - 18.43 19.98 30.96
GraphGPT - - 70.11 - - - - - - - -
UniGraph-Llama-7B 69.53 89.74 72.48 38.45 66.07 75.73 43.45 60.23 36.73 38.24 54.32
UniGraph-Llama2-7B 66.39 86.35 74.45 39.23 68.98 76.11 44.21 58.45 37.86 40.21 55.19
UniGraph-Llama2-13B 70.24 89.99 74.98 40.37 70.34 78.44 45.90 65.65 38.21 42.01 56.12
  • 1
Table 15. Results of zero-shot transfer compared to ChatGPT. We report accuracy (%) for all datasets.
Cora PubMed Products1 Wiki-CS WN18RR
GPT-3.5 Turbo 67.69 - 93.42 74.40 - - - - - - -
UniGraph-Llama-7B 69.53 89.74 72.48 38.45 66.07 75.73 43.45 60.23 36.73 38.24 54.32
UniGraph-Llama2-7B 66.39 86.35 74.45 39.23 68.98 76.11 44.21 58.45 37.86 40.21 55.19
UniGraph-Llama2-13B 70.24 89.99 74.98 40.37 70.34 78.44 45.90 65.65 38.21 42.01 56.12
  • 1
@@ Appendix D Graph Neural Networks (GNNs)
Appendix E Extended Related Work
@@
Single graph learning. Graph Neural Networks(GNNs) (Kipf and Welling, 2016a; Veličković et al., 2018a) take node features as input and aggregate local neighbor representations using the message-passing paradigm, directly optimizing for specific downstream tasks, achieving superior performance. With GNNs as the backbone, graph self-supervised learning (Kipf and Welling, 2016b; Veličković et al., 2018b) learns representation extractors on unlabeled graphs, subsequently applying the representations to downstream tasks or fine-tuning the pretrained GNNs. Unlike NLP tasks, graph learning tasks exhibit considerable diversity in their forms, making it a challenge to adapt a graph model to different downstream tasks. Recently, graph prompt learning has attempted to unify all tasks into either edge-level (Liu et al., 2023b) or graph-level (Sun et al., 2023) tasks.
@@
Large language models on graphs. Many real-world graphs naturally come with text as node or edge features, which we refer to as Text-Attributed Graphs(TAGs). To facilitate the learning of graph models, using language models to encode text into low-dimensional vectors is a common practice (Hu et al., 2020). Furthermore, recent Large Language Models(LLMs), represented by ChatGPT, demonstrate extensive common knowledge and capabilities as general task solvers. Some works transform graph structures into text and combine them with task descriptions to form prompts as input for LLMs, attempting to directly use LLMs as predictors for handling graph learning tasks (Guo et al., 2023; Wang et al., 2023; Liu and Wu, 2023). Furthermore, some research attempts to enhance the ability of LLMs to understand graph structures and tasks by employing instruction tuning (Ye et al., 2023; Tang et al., 2023; Xiong et al., 2024; Yang et al., 2023; Sui et al., 2024).
Technical contributions of UniGraph. Existing graph learning models on TAGs are either GNN-nested Transformers like GraphFormers (Yang et al., 2021) or models like GLEM (Zhao et al., 2022) and TAPE (He et al., 2024a) that decouple LMs and GNNs, optimizing them separately. The primary reasons are due to the large parameter size of LMs and the connectivity of graph structures making the model difficult to scale. Additionally, simultaneously optimizing LMs and GNNs with a significant disparity in parameter size can lead to phenomena such as overfitting. Papers like GLEM (Zhao et al., 2022) have mentioned the difficulties in training such a model. In UniGraph, we utilized subgraph sampling and a self-supervised learning loss function based on MLM to effectively mitigate these issues, making this type of backbone network a viable option.
Appendix F Connections to Graph Out-Of-Distribution (OOD) Generalization
While they share a common ground in aiming to enhance the model’s generalization capabilities, they have distinct objectives, methodologies, and applications.
F.1. Similarities
Both approaches aim to improve the model’s generalization capabilities. Our model seeks to create a robust, pre-trained model that can be adapted to various specific tasks or domains. Similarly, Graph OOD Generalization focuses on ensuring that models perform well on graph data that follow a different distribution from the training set.
F.2. Differences
Objectives: UniGraph, while capable of generalizing across domains, primarily focuses on leveraging large-scale data for comprehensive self-supervised pre-training, with less emphasis on specific distributional shifts. Graph OOD Generalization places a strong emphasis on the model’s ability to handle data from distributions that are not represented in the training set, usually supervised learning.
Model architecture and training: UniGraph is a large, pre-trained model that learns from vast amounts of graph data, capturing a wide variety of patterns and features. The training process involves self-supervised pre-training. In contrast, Graph OOD Generalization strategies may be integrated into various model architectures without the necessity for a large-scale pre-training phase. Techniques such as domain adaptation, invariant feature learning, or meta-learning are often employed to improve OOD performance.
Application scope: UniGraph aims to serve as a versatile foundation for a wide range of downstream tasks and domains by leveraging the power of pre-training. As discussed in our paper, the difficulty lies in the fact that graphs from different domains have different features and label spaces; we choose to use text to unify these spaces. Existing methods for graph OOD (Out-Of-Domain) generalization often focus on the same domain and the same feature/label dimensions, where the training and test sets have inconsistent graph structures and feature distributions (Li et al., 2022; Zhang et al., 2021c). This is primarily represented by transfer learning tasks on molecular graphs (Li et al., 2022). To the best of our knowledge, there is currently no paper focused on addressing the issue of cross-domain graph feature/label space misalignment.
Appendix G Instruction Prompts
@@
Knowledge graphs. Given a knowledge graph, edge between two entities represents a relation with a specific category. Node one <$\text{node}_{v}$> is featured with its content: [Name],[Content]. Node two <$\text{node}_{u}$> is featured with its content: [Name],[Content]. Question: Which category should the relation between node one Node one <$\text{node}_{v}$> and node two <$\text{node}_{u}$> be classified as? Please strictly classify the Wikipedia page into one of the following categories:[Candidate Labels]. Answer:
Appendix H Analysis of LLMs Choices
In Table 14, we report zero-shot transfer results with different LLMs.
Appendix I Results of Zero-Shot Transfer Compared to ChatGPT.
In table 15, ChatGPT(gpt-3.5-turbo) achieves better performance on most datasets, which we speculate is due to its large number of parameters and extensive pre-training data.
Appendix J On Incorporating Additional Graph Zero-Shot Learning Baselines
Table 16. Zero-shot results comparison between UniGraph and OpenGraph on Cora and PubMed datasets.
Method Cora PubMed
OpenGraph 75.04 68.69
UniGraph 78.24 74.21
Table 16 compares the zero-shot results of UniGraph and OpenGraph (Xia et al., 2024) on the Cora and PubMed datasets. We adopt an experimental setup strictly consistent with OpenGraph, pre-training on the pre-training datasets provided by OpenGraph.
Table 17. Zero-shot results comparison between UniGraph and LLaGA on Arxiv, Products, Cora and PubMed datasets.
Method Arxiv Products Cora PubMed
LLaGA 74.29 82.21 92.42 87.82
UniGraph 76.01 82.41 93.12 87.98
Table 17 compares the zero-shot results of UniGraph and LLaGA (Chen et al., 2024) on the Arxiv, Products, Cora and PubMed datasets. We adopt an experimental setup strictly consistent with LLaGA, pre-training on the pre-training datasets provided by LLaGA.
Table 18. Zero-shot results comparison between UniGraph and ZeroG on Cora and PubMed datasets.
Method Cora PubMed
ZeroG 68.72 78.02
UniGraph 72.19 81.23
Table 18 compares the zero-shot results of UniGraph and ZeroG (Li et al., 2024) on the Cora and PubMed datasets. We adopt an experimental setup strictly consistent with ZeroG, pre-training on the pre-training datasets provided by ZeroG.

### v3

@@ Abstract.
Graph Neural Networks; Self-Supervised Learning; Graph Pre-Training
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709277isbn: 979-8-4007-1245-6/25/08ccs: Information systems Data miningccs: Computing methodologies Neural networksccs: Information systems Social networks
@@ 4.2. Graph Siamese Masked Autoencoders on TAGs
The existing graph self-supervised learning methods (Veličković et al., 2018b; Kipf and Welling, 2016b; Thakoor et al., 2021; Hou et al., 2022) adopt GNNs as the backbone networks and use pre-processed vector features as input. In this part, we propose an end-to-end self-supervised learning method on TAGs. We cascade a pre-trained language model (LM) and a GNN as the backbone network. In this paper, we adopt DeBERTa-base (He et al., 2020) as our LM and GAT (Veličković et al., 2018a) as our GNN. Inspired by successful self-supervised learning techniques such as masked modeling (Kenton and Toutanova, 2019; He et al., 2022; Hou et al., 2022; Zhao et al., 2024) and siamese networks (Wu et al., 2018), we designed Graph Siamese Masked Autoencoders, enabling large-scale self-supervised pre-training on TAGs.
@@
Table 1. Statistics of all eleven text-attributed graph datasets.
Dataset Domain Task #Graphs Avg. #Nodes Avg. #Edges Raw Texts
Cora Citation Node 1 2,708 5,429 paper titles and abstracts
PubMed Citation Node 1 19,717 44,338 paper titles and abstracts
ogbn-Arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
ogbn-Papers100M Citation Node 1 111,059,956 1,615,685,872 paper titles and abstracts
ogbn-Products Product Node 1 2,449,029 61,859,140 product descriptions
Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
FB15K237 Knowledge Edge 1 14,541 310,116 entity names and descriptions
WN18RR Knowledge Edge 1 40,943 93,003 entity names and descriptions
PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
ChEMBL Molecule Graph 365,065 25.9 55.9 textual descriptions of atoms/bonds
@@
    Due to the difference between the pre-trained datasets used in their original paper and ours, we re-run the experiments using their open-source code.
4.3. Graph Instruction Tuning
Instruction prompts. For instruction tuning, we design prompt templates that include graph embeddings, graph structure, and natural language instructions, as summarized in Appendix D.
5. Experiments
5.1. Self-Supervised Representation Learning
Setup. We adopt the most commonly used linear probing protocol to evaluate the representation learning ability of self-supervised pre-trained models on unseen datasets. We train a linear classifier on top of the embeddings from a frozen model. Our model and all self-supervised learning baseline methods are first pre-trained on the large-scale citation network ogbn-Papers100M (Hu et al., 2020). Then, we evaluate on nine graphs from five domains with different tasks. Detailed settings and hyper-parameters can be found in Appendix B.
5.2. Few-Shot Transfer
5.3. Zero-Shot Transfer
Setup. For zero-shot transfer, we mainly compare with general-purpose LLMs such as Llama (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) and vicuna-7B-v1.5 (Chiang et al., 2023), as well as graph-based LLM methods GraphGPT (Tang et al., 2023), ZeroG (Li et al., 2024) and G2P2 (Wen and Fang, 2023). Also, we compare with OFA mentioned in section 5.2. Consistent with OFA, we select ogbn-Arxiv and FB15K237 as instruction tuning datasets. Note that there is no overlap between their label categories and those of the downstream datasets. We report the average accuracy on the original test set with 3 random initializations for each downstream dataset.
@@ 6. Conclusion
Acknowledgements.
This research is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 2 (FY2025) (Award MOE-T2EP20124-0009).
@@ Appendix A Datasets
Appendix B Implementation Notes
@@
Appendix C Extended Related Work
@@
Single graph learning. Graph Neural Networks(GNNs) (Kipf and Welling, 2016a; Veličković et al., 2018a; Yang et al., 2024) take node features as input and aggregate local neighbor representations using the message-passing paradigm, directly optimizing for specific downstream tasks, achieving superior performance. With GNNs as the backbone, graph self-supervised learning (Kipf and Welling, 2016b; Veličković et al., 2018b) learns representation extractors on unlabeled graphs, subsequently applying the representations to downstream tasks or fine-tuning the pretrained GNNs. Unlike NLP tasks, graph learning tasks exhibit considerable diversity in their forms, making it a challenge to adapt a graph model to different downstream tasks. Recently, graph prompt learning has attempted to unify all tasks into either edge-level (Liu et al., 2023b) or graph-level (Sun et al., 2023) tasks.
@@
Large language models on graphs. Many real-world graphs naturally come with text as node or edge features, which we refer to as Text-Attributed Graphs(TAGs). To facilitate the learning of graph models, using language models to encode text into low-dimensional vectors is a common practice (Hu et al., 2020). Furthermore, recent Large Language Models(LLMs), represented by ChatGPT, demonstrate extensive common knowledge and capabilities as general task solvers. Some works transform graph structures into text and combine them with task descriptions to form prompts as input for LLMs, attempting to directly use LLMs as predictors for handling graph learning tasks (Guo et al., 2023; Wang et al., 2023; Liu and Wu, 2023; Sui et al., 2024a). Furthermore, some research attempts to enhance the ability of LLMs to understand graph structures and tasks by employing instruction tuning (Ye et al., 2023; Tang et al., 2023; Xiong et al., 2024; Yang et al., 2023; Sui et al., 2024b).
Appendix D Instruction Prompts
@@
Knowledge graphs. Given a knowledge graph, edge between two entities represents a relation with a specific category. Node one <$\text{node}_{v}$> is featured with its content: [Name],[Content]. Node two <$\text{node}_{u}$> is featured with its content: [Name],[Content]. Question: Which category should the relation between node one Node one <$\text{node}_{v}$> and node two <$\text{node}_{u}$> be classified as? Please strictly classify the relation into one of the following categories:[Candidate Labels]. Answer: